y=6+x-x^2
find vertex, determine if it is max or min point, find zeros if exists
vertex: x=-b/2a
= (1)/((2(-1))= 1/-2
f(1/-2)=6+ (1/-2)^2=6.25 since a=1 parabola goes down x=-2 x=3
I am not sure if i did this right?

Answer by ewatrrr(24785)

y=-x^2 + x + 6 = -(x-.5)^2 + .25 + 6 = -(x-.5)^2 + 6.25
V(.5, 6.25) max (Parabola Opening downward)
-(x-.5)^2 + 6.25 = 0
x - .5 = ± 2.5
x = .5 ± 2.5
Xeros
x = -2, 3